OpenStudy (anonymous):

Given the function f(x) = −2x + 4, find x if f(x) = 20. −8 −36 44 −12

OpenStudy (anonymous):

do the same as the previous problem... what have you tried?

OpenStudy (anonymous):

where did you get stuck?

OpenStudy (anonymous):

-2x+4=20 -2x=16 divide by -2 so x=-8